Consider the die-toss experiment. Define the following events:
A: {Toss an even number}
B: {Toss a number less than or equal to 3}
a) Describe
1
Expert's answer
2020-08-04T19:46:46-0400


A die has 6 sides.


Let X be the outcome of a die-toss


"X =\{1,2,3,4,5,6\}"


Solution A: Toss an even number

"X \, is \, even = \{2,4,6\}"


Solution B: Toss a number less than or equal to 3

"X \leq 3 = \{1,2,3\}"


Solution a) Describe


Assuming a fair die


"p(X = x) = \frac{1}{6}"


"p(X \, is\, even) = p(X =2)+ p(X=4)+ p(X=6)"

"= \frac{1}{6}+\frac{1}{6}+\frac{1}{6} = \frac{1}{2}"


"p(X \leq 3)=p(X=1)+ p(X=2)+ p(X=3)"


"=\frac{1}{6}+\frac{1}{6}+\frac{1}{6}=\frac{1}{2}"